OpenStudy (anonymous):

Please help!? I do not understand how to do this at all. Volume of HCl solution 100.5 mL HCl Volume of NaOH solution 100.0 mL NaOH Initial temperature in calorimeter 25.2 °C Final temperature in calorimeter 28.2 °C 3. Determine the number of moles of NaOH. 4. Determine enthalpy per mole of NaOH. Show all of your work. enthalpy change of this reaction= 1.258 kj/mol

OpenStudy (anonymous):

Are we to assume that there was no reagent in excess?

OpenStudy (anonymous):

Assuming there is no reagent in excess, Find the energy change of the system using \[\Delta Q = mc \theta\] m = mass of liquid in the beaker (assuming they're diluted, 1 mL of liquid = 1g) c = specific heat capacity of water theta = temperature change (initial temp - final temp) (Note the energy change, the enthalpy change of the reaction and temperature change are all negative values, so enthalpy change of this reaction = -1.258 kJ/mol) The enthalpy change of this reaction would be the enthalpy change of neutralisation \[\Delta H [neutralisation] = \frac{ \Delta Q }{ n(water) }\] n(water) = moles of water Note: Convert the enthalpy change of neutralisation to J/mol first You can then find the moles of water, and using stoichiometry, you can find the moles of NaOH. The enthalpy per mole of NaOH would be \[\Delta H [NaOH] = \frac{ \Delta Q }{ n (NaOH) }\] Whicih is the same as the enthalpy change of the reaction